OpenStudy (anonymous):

find an equation of the form y-k=a(x-h)^2 for the parabola that has a vertex (1,-5) and contains (-4,3)

OpenStudy (anonymous):

Find an equation of the form y-k = a(x-h)[squared] for the parabola that has vertex (1, -5) and contains (-4, 3). --------------------- Vertex is (h,k) So, h = 1 and k = -5 ------ Form: y+5 = a(x-1)^2 ----- Using (-4,3) you can solve for "a": 3+5 = a(-4-1)^2 a = 8/25 ---- Equation: y + 5 = (8/25)(x-1)^2 I hope this helps!!

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

No problem!